1. 2x^2 - 7x + 5 =0
2. 3x^2 + 10x + 5 =0
3. -5 + or - the square root of 10 / 3


Answer by ikdeep(226)

Solution 1.
2x^2 - 7x + 5 =0
Here you must know how we make factors i.e. splitting of the middle term...if you might not knew then contact me again and I will explain to you)...
2x^2 - (5+2)x + 5 =0
opening brackets would give ,,,
2x^2 - 5x -2x + 5 =0
taking common from first two i.e. 2x^2 - 5x and last two i.e. -2x + 5 variables...we get

2x(x-5) -1(x-5) = 0 ...
(2x-1)(x-5)= 0 ..............ANSWER
hope this will help you
Please feel free to revert back for any further queries.